There are three Groups A,B and C with no. of persons 50, X, 90. The mean and Standard deviations for these groups are 113, Y, 115 and 6, 7, Z respectively. The Combind n , mean and Standard deviation for three groups are 200,116 and 7.746.




(i) What is n ( The value of X ) for Group II .




(ii) Mean (Value of Y ) of Group II?




(iii) S.D. (Value of Z) of Group III ?

1
Expert's answer
2022-02-23T10:33:29-0500

(i)


"50+X+90=200"

"X=60"

(ii)


"\dfrac{50(113)+60Y+90(115)}{200}=116"

"6Y=2320-565-1035"

"Y=120"

(iii)


"(7.746)^2=\dfrac{50(6)^2+60(7)^2+90(Z)^2}{200}"

"+\dfrac{50(113-116)^2+60(120-116)^2+90(115-116)^2}{200}"

"90(Z)^2=5760.1032"

"Z=8"